# Understanding intra-node communication in HPC systems and Datacenters

### v1

@@ Abstract
Over the past decade, specialized computing and storage devices, such as GPUs, TPUs, and high-speed storage, have been increasingly integrated into server nodes within Supercomputers and Data Centers. The advent of high-bandwidth memory (HBM) has facilitated a more compact design for these components, enabling multiple units to be interconnected within a single server node through intra-node networks like PCIe, NVLink, or Ethernet. These networks allow for scaling up the number of dedicated computing and storage devices per node. Additionally, inter-node networks link these devices across thousands of server nodes in large-scale computing systems. However, as communication demands among accelerators grow—especially in workloads like generative AI—both intra- and inter-node networks risk becoming critical bottlenecks. Although modern intra-node network architectures attempt to mitigate this issue by boosting bandwidth, we demonstrate in this paper that such an approach can inadvertently degrade inter-node communication. This occurs when high-bandwidth intra-node traffic interferes with incoming traffic from external nodes, leading to congestion. To evaluate this phenomenon, we analyze the communication behavior of realistic traffic patterns commonly found in generative AI applications. Using OMNeT++, we developed a general simulation model that captures both intra- and inter-node network interactions. Through extensive simulations, our findings reveal that increasing intra-node bandwidth and the number of accelerators per node can actually hinder overall inter-node communication performance rather than improve it.
Keywords High-performance interconnection networks  $\cdot$ Intra-node communication  $\cdot$ accelerators  $\cdot$ GPUs
@@ 1 Motivation
Supercomputers and Data centers are increasing their computing power and storage capacity to cope with the growing requirements of applications and services, such as those used in scientific computing or generative AI. While computing power requirements are currently being met using more powerful processors and specific-purpose accelerators (GPUs, TPUs, etc.), innovative technologies, such as 3D-stacked high-bandwidth memory (HBM), Non-Volatile Memory (NVMe) or Storage Class Memory (SCM), are increasing storage and memory capacity and bandwidth access, to bring unprecedented amounts of data close to where it is processed and reduce inter-node communication latency.
@@
The devices mentioned above (i.e., accelerators, CPUs, memory, or storage) are interconnected within end nodes through an intra-node interconnection network (also known as I/O or system interconnect), which must be designed so that it does not become the bottleneck in the communication among these devices. Generally, this is achieved by over-dimensioning the aggregated bandwidth of the intra-node network links. For instance, the Peripheral Component Interconnect Express (PCIe) is the de-facto intra-node network in contemporary servers, offering 1GB/s per lane for PCIe 3.0, 2GB/s for PCIe 4.0, and 4GB/s for PCIe 5.0. Other intra-node technologies exist such as NVLink (NVIDIA), which is specially designed to interconnect multiple GPUs at either the same end-node or different end-nodes, reaching speeds of 4GB/s, 7.5GB/s, and 15.1 GB/s per lane for Gen5, Gen6, and Gen7 (planned for next year), respectively. Infinity Fabric (AMD) is also used to interconnect specific AMD devices for both CPU (i.e., Zen) and graphics (e.g., Vega). Therefore, the increasing link speed in the intra-node network and the variety of devices within a server node (i.e., CPUs, GPUs, TPUs, other accelerators, SSDs, NVMe, etc.) challenges the design of intra-node network technologies, which needs to thoroughly identify potential bottlenecks when different communication patterns are generated involving multiple end-node devices that request to communicate through the intra-node network at full speed.
@@
In this context, communication operations among end-node devices (e.g., GPUs and other accelerators) may involve devices beyond a single node. For instance, an AllReduce communication operation performed when training a deep neural network (DNN) usually involves multiple GPUs that can be placed at different end nodes when the model to train does not fit a single end node [1]. Following this example, when the model is too big to be stored in the memory of a single end-node device, it is usually stored in the local memory of multiple end-node devices (e.g., the HBM local memory of GPUs). So, when the weights of that model need to be updated in a distributed manner, communication operations are generated involving multiple end-node devices, which will encapsulate the communication operations information in intra-node network transactions that will be exchanged among the devices involved in the communication, which may be placed at different end-nodes. More precisely, a communication operation involving two end-node devices (e.g., GPUs) can be divided into three phases. First, an end-node device generates an intra-node network transaction with the specific information to be transmitted. This transaction is divided into intra-node network packets and sent throughout the intra-node network to another device within the same end node or the network interface (NIC). In any of these cases, the receiving end-node device finally composes the transaction based on the received packets. Secondly, when the transaction is sent to a remote end-node device from the mentioned NIC, the intra-node transaction is encapsulated into inter-node network packets and sent through the inter-node network to the NIC at the destination end node. Third, at the NIC at the receiving end-node, packets are combined into an intra-node network transaction, divided into intra-node packets, and sent through the intra-node network reaching the destination end-node device, which finally processes that transaction. Note that the interconnection network performance has been widely studied in the last decades only assuming the second communication phase. However, the impact of communication phases first and second in the intra- and inter-node network performance is still a wide-open field of study [2]. Therefore, it is critical to accurately analyze the communication operations generated at end-node devices under specific communication patterns and their impact on the intra- and inter-node network performance. This analysis will help identify potential communication bottlenecks in the system.
Note that the aggregated communication bandwidth that multiple end-node devices may require at a given instant can also be requested to the inter-node interconnection network when devices at different end nodes want to communicate. Consequently, the inter-node network bandwidth is augmenting proportionally to accommodate the communication requirements of the increasing number of devices per end node and their computing power and storage requirements. Furthermore, current inter-node interconnection network technologies are focused on proposing more cost-effective network topologies, routing algorithms, and specific mechanisms to improve network performance in several ways, such as congestion control, power management, QoS, network virtualization, fault tolerance, etc.
@@
This paper addresses the following research question: to what extent does the intra-node network communication impact the inter-node communication and vice-versa? To answer this question, we have characterized the intra-node network architecture and the communication operations (intra- and inter-node) generated at end-node devices and modeled them into an OMNeT++-based simulation tool. Note that there are several proposals for system-level simulators (e.g., gem5 [3]), which model accurately the processor, memory, GPUs, and I/O. Other solutions model the specific GPU-to-GPU communication (e.g., TraceR-CODES [4, 5]). However, these models lack scalability or an accurate performance model at the packet level for the combined intra-node and inter-node communication. We have validated our simulation model using a real cluster infrastructure. We have run specific micro-benchmarks (e.g., InfiniBand Perftest and OSU benchmarks) and obtained latency and throughput metrics, which we have compared to those obtained using our simulation model. Moreover, we have performed simulation experiments configuring intra- and inter-node networks that communicate $32$ and $128$ end-nodes (and $256$ and $1024$ accelerators, respectively) using realistic communication patterns, such as those used in the training of Large Language Models (LLMs) for generative AI. The obtained results show that the intra-node network can be a bottleneck, due to the interference between intra-node and inter-node communication operations, increasing dramatically the tail latency at high intra- or inter-node traffic loads.
@@
The rest of the paper is organized as follows. Section 2 overviews the background of end-node devices, inter- and intra-node interconnects, and communication patterns. Section3 describes our proposed simulation model to analyze the impact of intra- and inter-node communication interference. Section 4 describes the simulation experiments and discusses the obtained results. Finally, Section 5 draws some conclusions.
@@ 2 Background
This section provides background details on end-node devices and the interconnection network architecture. Moreover, we analyze critical communication patterns from typical applications currently run in supercomputers and data centers.
@@ 2.1 End-node devices
The Central Processing Unit (CPU) is the central component at server nodes in supercomputers or data centers. The CPU manages all processes and jobs within the node and interprets and executes instructions to coordinate software and hardware operations. For instance, in generative AI applications based on Large Language Models (LLMs), the CPU prepares and preprocesses vast datasets before sharing them to GPUs or TPUs for training.
The Graphics Processing Unit (GPU) is a specialized processor optimized for computations requiring massive parallelism. Unlike CPUs, which excel in sequential task execution, GPUs feature thousands of smaller cores designed for simultaneous operations. This architecture makes GPUs critical for the training and inference in LLMs. Leading manufacturers such as NVIDIA, AMD, and Intel have developed GPUs tailored for intensive AI workloads. For instance, the NVIDIA Hopper GPU [6] is one of the most advanced GPUs of this company. It combines high memory bandwidth (HBM), increased computational power, and seamless communication via the specific NVLink intra-node network. The NVIDIA Hopper GPU is integrated into the Grace Hopper Superchip [7], which combines the NVIDIA Hopper GPU with the NVIDIA Grace CPU. It delivers up to $989.4$ TFLOPS in FP32 and features up to $144$GB of HBM3e memory, making it a famous solution for AI training and inference. Alternatively, the Intel Gaudi GPU [8] is a powerful GPU designed for large-scale AI training and inference. It provides support for massive datasets and seamless integration into various systems. Intel Gaudi GPUs deliver up to $229$ TFLOPS with $128$GB of HBM memory, making it a serious contender in the AI hardware market.
Solid State Drives (SSDs) and other high-speed storage technologies (e.g., NVMe or SCM) are also key components in data centers and supercomputers. They provide the necessary storage capacity and data retrieval speeds to support the large datasets typical of generative AI and HPC workloads. The performance of memory disks directly impacts data transfer rates, latency, and overall system efficiency. Emerging technologies, such as NVMe over Fabrics (NVMe-oF) enhance storage performance, bridging the gap between compute and data requirements in large-scale systems.
The end-node devices mentioned above are interconnected using an intra-node network. They can also communicate with other devices placed at different end nodes using an inter-node network. The network interface (NIC) processes intra-node packets and injects them into the inter-node network.
@@ 2.2 Inter-node interconnection networks
They are a critical subsystem in Supercomputers and Datacenters to provide high-performance communication operations among thousands of end nodes at a reduced cost; reliability, since large-scale interconnection networks increase fault probability; and energy efficiency to minimize the network energy fraction.
@@
Numerous manufacturers provide fast and efficient network technologies, such as NVIDIA and its InfiniBand-based network devices, HPE and Cray, which promote the Slingshot technology, or all the partners composing the recently created Ultra Ethernet Consortium (UEC), whose primary goal is to optimize the Ethernet network technology for high-performance AI and HPC. For instance, the RDMA over Converged Ethernet (RoCE) is an advanced protocol implemented in NVIDIA devices that enables Remote Direct Memory Access (RDMA) over Ethernet networks. RoCE is designed to reduce CPU overhead, lower latency, and improve throughput by bypassing traditional TCP/IP stacks. It is widely used in inter-node communication within Datacenters, especially for storage and high-performance computing (HPC). An interesting implementation of RoCE can be found in Intel’s Gaudi3 system, which leverages RoCE to optimize the performance of AI training workloads. By integrating RDMA capabilities directly into the system architecture, Gaudi3 achieves efficient inter-node communication while minimizing latency and maximizing data throughput.
@@ 2.3 Intra-node interconnection networks
They are also a fundamental subsystem at end nodes of Supercomputers and Data centers. Intra-node networks must offer ultra-low latency to facilitate communication between CPUs, accelerators, and memory; high bandwidth to support data-intensive workloads; and energy efficiency to ensure sustainable performance. Furthermore, as end nodes become increasingly heterogeneous due to the combination of CPUs, GPUs, and specialized accelerators, the intra-node networks cope with new communication patterns. Historically, these patterns have been simple, so the intra-node networks were based on bus-based or ring-based interconnects. However, modern intra-node networks have evolved significantly to accommodate the demand for higher performance and scalability. Nowadays, intra-node networks support emerging workloads, such as generative AI models that demand unprecedented levels of parallelism and data movement efficiency [11]. Consequently, intra-node interconnection networks incorporate innovations such as in-network hardware accelerators for specific communication patterns, advanced routing algorithms, novel flow control techniques, power-aware communication strategies, etc.
PCIe is the mainstream interconnect between CPUs, GPUs, network cards, and storage devices. The PCIe bandwidth and low latency enable efficient data transfers crucial for modern high-performance workloads. PCIe speeds have evolved significantly, with PCIe 5.0 and PCIe 6.0 offering per-lane speeds of up to 3.9 and 7.5 GB/s, respectively. These advancements improve the throughput and reduce communication bottlenecks. PCIe is used, for instance, to scale up the Intel’s HL-325L Habana Gaudi 3 accelerators within a single node. Furthermore, PCIe is a foundational interconnect for inter-node network interface cards (NICs), directly impacting the overall system performance by reducing data movement delays and enhancing throughput. By contrast, NVLink (NVIDIA) provides significantly higher bandwidth than PCIe, with the latest versions supporting up to 900 GB/s of bidirectional bandwidth. This capability enables seamless data sharing between GPUs, CPUs, and memory subsystems, making it particularly effective for AI/ML workloads and other compute-intensive applications. NVLink is also employed in certain inter-node contexts, such as NVIDIA DGX systems, where it enhances the scalability and performance of multi-node configurations by creating tightly coupled GPU clusters.
There are other intra-node network solutions, such as Infinity Fabric (AMD) or Quick Path Interconnect (Intel), and even open-standard efforts such as the UALink Consortium that are also playing an important role in the intra-node networks design. Furthermore, emerging standards and protocols, such as the Compute Express Link (CXL), aim to unify memory and compute interconnects, enhancing intra-node communication efficiency.
@@ 2.4 Communication Patterns
The communication operations generated by emerging applications may threaten the performance of intra- and inter-node interconnection networks. With the rapid growth of generative AI applications, dominated by Large Language Models (LLMs), the size of these models are expanding so they no longer fit on a single GPU or even in the GPUs of a single end node. Therefore, parallelization is needed to accelerate LLM training and inference. The parallel programming model used to leverage the computing power and storage capacity of these systems is currently based on two main types of parallelism: data parallelism and model parallelism.
On the one hand, Data Parallelism (DP) [12] distributes the training process across multiple computational devices by dividing the dataset into smaller chunks (or batches). Each device (or group of devices) processes its assigned data portion independently, using an identical model copy. DP enables efficient scaling and faster processing. During training, communication between model copies occurs after a batch is processed when gradients are computed and updated. At this stage, each model copy exchanges gradient updates with all other copies using an AllReduce communication operation. This communication step ensures consistent parameter updates across all model replicas.
@@
3 Intra- and inter-node simulation model
This section describes the intra- and inter-node interconnection network simulation model that we have developed to study the impact of intra-node communication on the performance of the intra- and inter-node interconnection network. We also describe the characterization of realistic communication patterns as those generated by emerging applications, such as LLMs. In the following sections, we describe the server node architecture that has inspired the generic simulation model, the characterization of the PCIe communication on that server node, the proposed generic intra- and inter-node simulation model, and the network traffic model that mimics the behavior of realistic communication operations in LLM parallel training.
3.1 Baseline end-node architecture
Since we want to validate the simulation model against a real infrastructure before performing scale-out simulations, we have used an end-node architecture based on PCIe 3.0 as the baseline for our simulation model. Figure 1 illustrates the architecture of an end node in the Cluster for the Evaluation of Low-Latency Architectures (CELLIA).

### v2

# On the Impact of Intra-node Communication in the Performance of Supercomputer and Data Center Interconnection Networks

@@ Abstract
In the last decade, specific-purpose computing and storage devices, such as GPUs, TPUs, or high-speed storage, have been incorporated into server nodes of Supercomputers and Data centers. The development of high-bandwidth memory (HBM) enabled a much more compact form factor for these devices, thus allowing the interconnection of several of them within a server node, typically using an intra-node interconnection network (e.g., PCIe, NVLink, or Infinity Fabric). These networks allow scaling up the number of specific computing and storage devices per node. Furthermore, the inter-node networks communicate thousands of these devices placed in different server nodes in a Supercomputer or Data Center. Unfortunately, the intra- and inter-node networks may become the system’s bottleneck due to the increasing communication demand among accelerators of applications such as generative AI. Although current intra-node network designs alleviate this bottleneck by increasing the bandwidth of the intra-node network, we show in this paper that such a high bandwidth for intra-node communication may hinder the inter-node communication performance when traffic from outside the node arrives at the intra-node devices, resulting in interference with intra-node traffic. To analyze the impact of this interference, we have studied the communication operations of realistic traffic patterns exploiting intra-node communication. We have developed a generic intra- and inter-node simulation model based on OMNeT++ and modeled the mentioned communication operations. We have also performed extensive simulation experiments that confirm that increasing the intra-node network bandwidth and the number of computing devices per node (i.e., accelerators) is counterproductive to the inter-node communication performance.
keywords:
High-performance interconnection networks, Intra-node communication, accelerators, and GPUs.
@@ 1 Motivation
Supercomputers and Data Centers are composed of thousands of server nodes, which are increasing their computing power, memory, and storage capacity to cope with the growing requirements of applications and services such as scientific computing, generative AI, or social networks. While computing power requirements are currently being met using more powerful processors and specific-purpose accelerators (GPUs, TPUs, etc.), innovative technologies, such as 3D-stacked high-bandwidth memory (HBM), Non-Volatile Memory (NVMe), or Storage Class Memory (SCM), are increasing the memory and storage capacity of server nodes and the bandwidth access to these devices. These technologies allow placing unprecedented amounts of data close to where they are processed, reducing inter-node communication latency since data processing is done at multiple accelerators plugged into the same server node.
@@
Accelerators and other devices are interconnected within server nodes through an intra-node interconnection network, which must be designed so that it does not become a bottleneck when these devices communicate. Typically, this is achieved by over-dimensioning the aggregated bandwidth of the intra-node network links. For instance, PCIe is the de facto intra-node network in contemporary servers, offering 1 GB/s per lane for PCIe 3.0, 2 GB/s for PCIe 4.0, and 4 GB/s for PCIe 5.0. As communication operations may involve devices beyond a single node, other intra-node technologies exist, such as NVLink (NVIDIA), which is specially designed to interconnect multiple accelerators at either the same or different server nodes (or end nodes). NVLink provides speeds of 4 GB/s, 7.5 GB/s, and 15.1 GB/s per lane for Gen5, Gen6, and Gen7, respectively. Infinity Fabric (AMD) is also used to interconnect specific AMD devices for both CPU (i.e., Zen) and graphics (e.g., Vega). Therefore, the increasing link speed in the intra-node network and the variety of devices within a server node challenge the design of intra-node network technologies.
This design needs to identify potential bottlenecks, assuming that millions of communication operations are generated in the intra- and inter-node networks involving thousands of devices at different server nodes. The most straightforward communication operation (i.e., peer-to-peer or P2P) involves two end-node devices (e.g., accelerators) and can be divided into three phases. First, the sender device generates a communication transaction with another device. This transaction is split into packets at the device’s intra-node network interface, which are sent throughout the intra-node network to another device within the same end node or to the network interface (NIC) connecting that end node to other end nodes. Secondly, if the transaction is addressed to a remote device, the intra-node packets are aggregated at the mentioned NIC into inter-node network packets, which are sent through the inter-node network to the destination end node. Third, at the destination end-node NIC, packets are split again into intra-node network packets and sent through the intra-node network to the destination device. Note that if a transaction is exchanged between devices within the same end node, the communication remains in the intra-node network of a single end node.
@@
By contrast, collective operations (e.g., AllReduce or Broadcast) used in parallel applications, such as the deep neural networks (DNNs) training, may involve multiple computing devices at different end nodes when the model to train does not fit into a single end node [1]. When this happens, the model is stored in a distributed manner in the local memory of multiple end-node devices (e.g., the local HBM memory of accelerators). The training process of DNNs requires updating the weights stored in the memory of the involved devices, so that collective communication operations encapsulate the weight information in network transactions exchanged among these devices, and may generate a high rate of traffic the intra- and inter-node network needs to handle efficiently. Although high-performance interconnection networks have been widely investigated in the last decades, the impact of communication phases first and third, mentioned above, in the intra- and inter-node network performance is still a wide-open field of study [2]. Indeed, it is critical to accurately analyze the network traffic patterns produced from communication operations at multiple end-node devices and their impact on the intra- and inter-node network performance. This analysis will help identify potential bottlenecks in the interconnection network.
@@
This paper addresses the following research question: To what extent does the inter-node network traffic impact the intra-node traffic and vice-versa? To answer this question, we have characterized a generic intra-node network architecture and the communication operations (intra- and inter-node) generated at multiple end-node devices. We have modeled the network architecture and generic but realistic traffic patterns into an OMNeT++-based simulation tool. Note that there are several proposals for system-level simulators (e.g., gem5 [3]), which accurately model the processor, memory, accelerators, and I/O. Other solutions model the specific accelerator-to-accelerator communication (e.g., TraceR-CODES [4, 5]). However, these models lack scalability or packet-level granularity for the combined intra-node and inter-node communication model. We have validated our simulation model using a real cluster infrastructure, where we ran specific micro-benchmarks (e.g., InfiniBand Perftest) and obtained latency and throughput metrics, which we have compared to those obtained using our simulation model. Moreover, we have performed simulation experiments configuring intra- and inter-node networks with different speeds that communicate $32$, $128$, and $512$ end nodes (with $256$, $1024$, and $4096$ accelerators, respectively) using realistic communication patterns, such as those used in the training of Large Language Models (LLMs). The scale-up and scale-out simulation analyses show that the intra-node network can be a bottleneck due to the interference between intra-node and inter-node communication and the overhead when aggregating intra-node packets into inter-node packets and splitting inter-node packets into intra-node packets. We also show that these issues can be alleviated, including specific features at network devices.
@@
The rest of the paper is organized as follows. Section 2 overviews the background of inter- and intra-node interconnects and communication patterns. Section 3 describes our proposed simulation model to analyze the impact of intra- and inter-node communication interference. Section 4 describes the simulation experiments and discusses the obtained results. Finally, Section 5 draws some conclusions.
@@ 2.1 End-node devices
The Central Processing Unit (CPU) at server nodes in supercomputers or data centers manages all processes or jobs and interprets and executes instructions to coordinate software and hardware operations. For instance, in generative AI applications based on Large Language Models (LLMs), the CPU prepares and preprocesses vast datasets before sharing them with GPUs or TPUs for training.
The Graphics Processing Unit (GPU) is a specialized processor optimized for computations requiring massive parallelism. Unlike CPUs, which excel in sequential task execution, GPUs and other accelerators feature thousands of smaller cores designed for simultaneous operations. This architecture makes GPUs critical for training and inference in LLMs. Leading manufacturers such as NVIDIA, AMD, and Intel have developed GPUs tailored for intensive AI workloads. For instance, the NVIDIA Hopper GPU [6] combines high memory bandwidth (HBM), increased computational power, and seamless communication via the specific NVLink intra-node network. The NVIDIA Hopper GPU is integrated into the Grace Hopper Superchip [7], which combines the NVIDIA Hopper GPU with the NVIDIA Grace CPU. It delivers up to $989.4$ TFLOPS in FP32 and features up to 144144GB of HBM3e memory, making it a famous solution for AI training and inference. Alternatively, the Intel Gaudi GPU [8] is a powerful GPU designed for large-scale AI training and inference. It supports massive datasets and seamless integration into various systems and delivers up to $229$ TFLOPS with 128128GB of HBM memory.
Solid State Drives (SSDs) and other high-speed storage technologies (e.g., NVMe or SCM) are also key components in data centers and supercomputers. They provide the necessary storage capacity and data retrieval speeds to support the large datasets typical of generative AI and HPC workloads. The performance of memory disks directly impacts data transfer rates, latency, and overall system efficiency. Emerging technologies, such as NVMe over Fabrics (NVMe-oF), enhance storage performance, bridging the gap between compute and data requirements in large-scale systems.
@@ 2.2 Inter-node interconnection networks
They are a critical subsystem in supercomputers and data centers to provide high-performance communication operations among thousands of end nodes at a reduced cost and reliability, since large-scale interconnection networks increase fault probability and energy efficiency to minimize the network energy fraction.
@@
Numerous manufacturers provide fast and efficient network technologies, such as NVIDIA and its InfiniBand-based network devices, HPE and Cray, which promote the Slingshot technology, or all the partners composing the recently created Ultra Ethernet Consortium (UEC), whose primary goal is to optimize the Ethernet network technology for high-performance AI and HPC. For instance, the RDMA over Converged Ethernet (RoCE) is an advanced protocol implemented in NVIDIA devices that enables Remote Direct Memory Access (RDMA) over Ethernet networks. RoCE is designed to reduce CPU overhead and latency, and to improve throughput by bypassing traditional TCP/IP stacks. It is widely used in inter-node communication within Data Centers, especially for storage and high-performance computing (HPC). An interesting implementation of RoCE can be found in Intel’s Gaudi3 system, which leverages RoCE to optimize the performance of AI training workloads. By integrating RDMA capabilities directly into the system architecture, Gaudi3 achieves efficient inter-node communication while minimizing latency and maximizing data throughput.
@@ 2.3 Intra-node interconnection networks
They are also a fundamental subsystem at the end nodes of Supercomputers and Data centers. Intra-node networks must offer ultra-low latency to facilitate communication between CPUs, accelerators, and memory; high bandwidth to support data-intensive workloads; and energy efficiency to ensure sustainable performance. Furthermore, as end nodes become increasingly heterogeneous due to the combination of CPUs, GPUs, and specialized accelerators, the intra-node networks cope with new communication patterns. Historically, these patterns have been simple, so the intra-node networks were based on bus-based or ring-based interconnects. However, modern intra-node networks have evolved significantly to accommodate the demand for higher performance and scalability. Nowadays, intra-node networks support emerging workloads, such as generative AI models that demand unprecedented levels of parallelism and data movement efficiency [11]. Consequently, intra-node interconnection networks incorporate innovations such as in-network hardware accelerators for specific communication patterns, advanced routing algorithms, novel flow control techniques, power-aware communication strategies, etc.
PCIe is the mainstream interconnect between CPUs, GPUs, network cards, and storage devices. The PCIe bandwidth and low latency enable efficient data transfers crucial for modern high-performance workloads. PCIe speeds have evolved significantly, with PCIe 5.0 and PCIe 6.0 offering per-lane speeds of up to 3.9 and 7.5 GB/s, respectively. These advancements improve the throughput and reduce communication bottlenecks. PCIe is used, for instance, to scale up Intel’s HL-325L Habana Gaudi 3 accelerators within a single node. Furthermore, PCIe is a foundational interconnect for inter-node network interface cards (NICs), directly impacting the overall system performance by reducing data movement delays and enhancing throughput. By contrast, NVLink (NVIDIA) provides significantly higher bandwidth than PCIe, with the latest versions supporting up to 900 GB/s of bidirectional bandwidth. This capability enables seamless data sharing between GPUs, CPUs, and memory subsystems, making it particularly effective for AI/ML workloads and other compute-intensive applications. NVLink is also employed in certain inter-node contexts, such as NVIDIA DGX systems, where it enhances the scalability and performance of multi-node configurations by creating tightly coupled GPU clusters.
Other intra-node network solutions, such as Infinity Fabric (AMD) or Quick Path Interconnect (Intel), and specifications such as the UALink, play an important role in the design of intra-node networks. Furthermore, emerging standards and protocols, such as the Compute Express Link (CXL), aim to unify memory and compute interconnects, enhancing intra-node communication efficiency.
@@ 2.4 Communication Patterns
The communication operations generated by emerging applications may threaten the performance of intra- and inter-node interconnection networks. With the rapid growth of generative AI applications, dominated by Large Language Models (LLMs), the size of these models is expanding so they no longer fit on a single accelerator or even in the accelerators of a single end node. Therefore, parallelization is needed to accelerate LLM training and inference. The parallel programming model used to leverage the computing power and storage capacity of these systems is currently based on two main types of parallelism: data parallelism and model parallelism.
Data Parallelism (DP) [12] distributes the training process across multiple computational devices by dividing the dataset into smaller chunks (or batches). Each device (or group of devices) processes its assigned data portion independently, using an identical model copy. DP enables efficient scaling and faster processing. During training, communication between model copies occurs after a batch is processed when gradients are computed and updated. At this stage, each model copy exchanges gradient updates with all other copies using an AllReduce communication operation. This communication step ensures consistent parameter updates across all model replicas.
@@
3 Modeling intra- and inter-node communication
This section describes the intra- and inter-node network simulation model developed to study the impact of intra-node communication on the performance of the intra- and inter-node interconnection networks, the characterization of the communication among PCIe devices, the generic intra- and inter-node simulation model, and the network traffic model reproducing realistic communication operations in Large Language Models (LLMs) parallel training.
3.1 Baseline end node architecture
In the simulation tool, we have modeled an end node architecture based on a real system using PCIe 3.0 to interconnect a GPU, an NVMe disk, and an InfiniBand NIC, which will be used for validation purposes. Figure 1 illustrates this infrastructure known as Cluster for the Evaluation of Low-Latency Architectures (CELLIA)  [^1] Further details on: https://www.i3a.uclm.es/raap/?page_id=1650..